Angle of elevation of top of a tower from two points at distances 10m and 5m from base of tower and in same straight line with it are complementary find height of tower

Grade:10

1 Answers

Abhijeet Pradhan
7 years ago
Tan a = h/5......Eqn1Tan (90-a) = h/10Cot a = h/10.......eqn2Multiplying eqn1 nd eqn2Tan a*cot a= h/10*h/51=(h^2)/50h^2=50h=√50=5√2m